# Adaptive Distribution Masked Autoencoders for Continual Test-Time Adaptation

### v1

@@ Abstract
Continual Test-Time Adaptation (CTTA) is proposed to migrate a source pre-trained model to continually changing target distributions, addressing real-world dynamism. Existing CTTA methods mainly rely on entropy minimization or teacher-student pseudo-labeling schemes for knowledge extraction in unlabeled target domains. However, dynamic data distributions cause miscalibrated predictions and noisy pseudo-labels in existing self-supervised learning methods, hindering the effective mitigation of error accumulation and catastrophic forgetting problems during the continual adaptation process. To tackle these issues, we propose a continual self-supervised method, Adaptive Distribution Masked Autoencoders (ADMA), which enhances the extraction of target domain knowledge while mitigating the accumulation of distribution shifts. Specifically, we propose a Distribution-aware Masking (DaM) mechanism to adaptively sample masked positions, followed by establishing consistency constraints between the masked target samples and the original target samples. Additionally, for masked tokens, we utilize an efficient decoder to reconstruct a hand-crafted feature descriptor (e.g., Histograms of Oriented Gradients), leveraging its invariant properties to boost task-relevant representations. Through conducting extensive experiments on four widely recognized benchmarks, our proposed method attains state-of-the-art performance in both classification and segmentation CTTA tasks.
@@ 1 Introduction
Deep Neural Networks (DNNs) have demonstrated impressive performance across various computer vision tasks, including image-level classification [18, 10, 31] and dense prediction [39, 55, 50], when the test data distribution closely aligns with the training data. However, this assumption is frequently challenged in real-world scenarios due to dynamic environments, with deployed models exhibiting insufficient generalization capabilities and performance degradation [42, 21]. Therefore, the problem of continual test-time adaptation (CTTA) has been introduced [47], aiming to adapt a source pre-trained model to continually changing target distributions. Due to privacy and practical considerations, during the continual adaptation process, access to source domain data is not permitted, and each target data can only be accessed once. While the CTTA showcases promising potential applications, it also increases the difficulty of transfer learning, which introduces catastrophic forgetting and error accumulation problems.
@@ 2.1 Continual Test-Time Adaptation
Test-time adaptation (TTA), also known as source-free domain adaptation [5, 27, 52], is the process of adapting a source model to an unknown target domain distribution without relying on any source domain data. Recent works have delved into techniques such as self-training and entropy regularization to fine-tune the source model [29, 6]. Tent [46] achieves this by updating the training parameters in batch normalization layers through entropy minimization. This approach has served as inspiration for subsequent research efforts in recent works [33, 54], which continue to investigate the robustness of normalization layers.
Continual Test-Time Adaptation (CTTA) denotes a scenario where the target domain is dynamically changing, introducing additional challenges for conventional TTA methods. The initial approach is presented in [47], which employs a teacher-student framework to integrate bi-average pseudo labels and stochastic weight reset. Drawing from the insight that mean teacher predictions are often more robust than standard models [44], a series of mainstream methods [13, 30, 9] continue this approach to self-supervised learning in CTTA. Specifically, [13] introduces the use of visual domain prompts to address the error accumulation and catastrophic forgetting problems at the input level. RMT [9] employs a symmetric cross-entropy loss to make the mean-teacher more robust. ViDA [30] introduces low-rank and high-rank visual domain adapters, explicitly handling both domain-specific and domain-shared knowledge. Concurrently, existing methods also focus on applying entropy minimization to update normalization layers [45, 17] or a fraction of model parameters [43]. EcoTTA introduces a meta-network to regularize the outputs from both the meta-network and the frozen network  [43]. However, due to continually changing environments, these self-training approaches are susceptible to miscalibrated predictions and noisy pseudo-labels, resulting in uncontrollable error accumulation. In this paper, we make the first attempt to explore the reconstruction mechanism in the CTTA problem and propose a novel approach to continual self-supervised learning, designed to enhance target domain knowledge extraction while mitigating distribution shift accumulation.
@@ 2.2 Masked Image Modeling
Mask-reconstruction-based self-supervised learning has been successful in reducing the reliance on extensive labeled datasets in both Natural Language Processing (NLP) and Computer Vision (CV). The concept was first introduced by BERT [8] in NLP, where a portion of the input word tokens is randomly masked, and the model learns to reconstruct the vocabularies of these masked tokens. In the field of computer vision (CV), similar techniques have been applied in various works [3, 20, 51]. These methods involve randomly masking a significant percentage of input image patches. Specifically, BEiT [3] was the first to explore Masked Image Modeling (MIM) in vision transformers by reconstructing the vision dictionary derived from DALL-E [37, 38]. MAE [20] scaled up MIM to larger models and demonstrated that a simple pixel reconstruction loss can enhance the visual representations of pre-trained models. However, relying solely on low-level RGB signals in MAE is considered rudimentary and limited in unlocking the full potential of MAE in downstream vision tasks. Subsequently, approaches like MaskFeat [48], data2vec [2], MVP [49], and MILAN [22] have uncovered various high-level signals, including pre-trained DINO features [34], hand-crafted features [7], momentum features [19], and multi-modality features [36]. Utilizing these high-level signals has been proven to be more effective in extracting contextual information. Furthermore, MR-MAE [16] adopts a unique approach by simultaneously learning both high-level and low-level signals during the pre-training.
@@ 3.1 Overview
Adaptive Distribution Masked Autoencoders One prevalent CTTA approaches primarily focus on applying entropy minimization to update the batch normalization layer [45, 17, 32] or a fraction of model parameters [43]. However, due to the dynamic environments, this self-training approach is susceptible to miscalibrated predictions, resulting in uncontrollable error accumulation. On the other hand, alternative mainstream approaches utilize the teacher-student pseudo-labeling in continual target domains. Nevertheless, the conventional mean teacher method [44] increases computational cost and produces noisy pseudo labels in dynamic environments, leading to the accumulation of distribution shifts. While methods such as [47, 30, 9] utilize Test-time augmentation methods to enhance the accuracy of pseudo labels, they may limit efficiency during the CTTA process. Different from prior continual self-supervised approaches, we lead the way in incorporating a masked autoencoder (MAE) to tackle the CTTA problem, abandoning the impact of miscalibrated predictions and the cumbersome teacher-student model. Our key insight lies in adopting the reconstruction scheme to effectively extract target domain knowledge while mitigating the accumulation of domain shift. The overall framework is illustrated in Figure 2. Specifically, We propose a Distribution-aware Masking (DaM) mechanism, detailed in Sec. 3.2, which adaptively places learnable masks on token embeddings with substantial distribution shifts. By establishing consistency constraints between the masked input and the original input, DaM significantly enhances the understanding of target domain knowledge and mitigates the challenge of error accumulation. For masked tokens, we adopt a linear decoder to reconstruct Histograms of Oriented Gradients, leveraging its invariant properties to acquire task-relevant representations while averting the introduction of target domain shifts, as elaborated in Sec. 3.3. This reconstruction method serves as a preventive measure, avoiding catastrophic forgetting in continual adaptation. We provide the intuitive explanation and justification in Sec. 5.
@@ 3.2 Distribution-aware Masking
To this end, as shown in Figure 2 (a), we introduce a Distribution-aware Masking (DaM) strategy, enabling a thoughtful selection of tokens to mask in dynamic environments. The key concept involves choosing tokens with substantial domain shifts for masking, ensuring that the preserved visible tokens exhibit relatively fewer domain shifts while providing reliable semantic knowledge through the model encoder. To quantify token-wise distribution shifts, we draw inspiration from [35, 40] and introduce a method for token-wise uncertainty estimation. Specifically, we employ the MC Dropout [11], enabling multiple forward propagations to obtain $m$ (e.g., $m=10$) sets of features for each token. Subsequently, we calculate the uncertainty value $\mathcal{U}(x)$ for a given token $x_{j}$, as formulated below:
@@ 4 Experiments
4.1 Experiments Details
Baseline. We compare our method with prior CTTA approaches, encompassing Pseudo-labeling [28], TENT [46], CoTTA [47], VDP [13], SVDP [53], and ViDA [30].
Implementation Details. In our CTTA experiments, to ensure consistency and fair comparisons, we follow the implementation details outlined in prior works [47, 30]. For the classification CTTA tasks, we employ the ViT-base [10] as our backbone model. We resize the input images to 384$\times$384 for CIFAR10-to-CIFAR10C, CIFAR100-to-CIFAR100C benchmark, and 224$\times$224 resolution for ImageNet-to-ImageNet-C benchmark. In the case of segmentation CTTA task, we employ the Segformer-B5 [50] pre-trained on the Cityscapes dataset as our source model. We down-sample the original input images resolution from 1920$\times$1080 to 960$\times$540. We use Adam [24] with $(\beta_{1},\beta_{2})=(0.9,0.999)$ as the optimizer. Different learning rates are assigned for different CTTA tasks and backbone models, such as we use 1e-5 for ViT on classification CTTA tasks and 3e-4 for Segformer on segmentation CTTA tasks. In the masking process, following the approach in MIC [23], we use [MASK] tokens initialized with all zeros to replace 50% of tokens that are embedded as patches. These [MASK] tokens are shared learnable embeddings that indicate masked patches. The reconstruction decoder is a randomly initialized linear layer used to project the output tokens associated with masked patches to HOG features.
@@ 4.1 Experiments Details
4.2 Classification CTTA Tasks
CIFAR10-to-CIFAR10C & CIFAR100-to-CIFAR100C. The source model is trained through supervised learning on the CIFAR10 or CIFAR100 dataset. During testing, we apply CTTA to the CIFAR10C or CIFAR100C dataset, which contains fifteen corruption types continuously fed into the model in a specific order. In the CIFAR10-to-CIFAR10C scenario, as shown in Table 1, the average classification error of the source model reaches 28.2% when directly testing on CIFAR10C. However, our method significantly reduces the error to 12.6%. Compared to other continual self-supervised methods, our approach outperforms them by 12.9% and 8.1% compared to the previous entropy minimization method (TENT [46]) and the state-of-the-art (SOTA) teacher-student method (ViDA [30]), demonstrating significant potential in addressing the CTTA problem. To be mentioned, our Adaptive Distribution Masked Autoencoders (ADMA) demonstrate outstanding performance across 12 out of the 15 corruption types, validating the robustness of our method in the continual adaptation process.
We expand our evaluation to the CIFAR100-to-CIFAR100C scenario, as depicted in Table 2, encompassing a more extensive range of categories within each domain. Our approach outperforms the previous entropy minimization and teacher-student pseudo-labeling methods by 6.8% and 8.4%, respectively, exhibiting a consistent trend with the aforementioned CTTA experiments. Therefore, the results validate the universality of our method, unaffected by the number of categories, and it can efficiently mitigate error accumulation and catastrophic forgetting problems.
@@ 4.4 Ablation Study
Error rate 49.2 48.5 48.1 42.5
Table 6: Average error rate(%) for ImageNet-to-ImageNet-C online CTTA task. RGB, CLIP, Mean-Teacher, and HOG are different reconstruction target features.
@@
Reconstruction target selection. Another set of ablation experiments aims to assess the impact of the reconstruction target, including the original RGB pixel, the high-level CLIP feature, the Mean-Teacher feature, and the HOG feature introduced in our method. Since the CIFAR datasets have an input size of $32\times 32$ with limited RGB pixel information, these ablation experiments are conducted in the context of the ImageNet-to-ImageNet-C. As shown in Table 6, our introduced HOG reconstruction outperforms other reconstruction targets by a significant margin. Specifically, compared with classical image pixel reconstruction, our method achieves a 6.7% improvement, demonstrating that reconstructing the RGB signal of the target domain introduces inherent domain shifts. In contrast to reconstructing the CLIP feature, our method outperforms it by 6.0%, while also significantly reducing computational costs as it does not require an additional model. The results validate that features directly extracted by the large-scale model still introduce domain shifts. Lastly, while reconstructing features of the Mean-Teacher model is slightly better (by 0.4%) than reconstructing CLIP features, it is still 5.6% lower than our method. In conclusion, reconstructing HOG and leveraging its invariant properties can boost task-relevant representations and avoid domain shift accumulation in CTTA.
@@ 5 Discussion and Justification
Inter-Domain Divergence. To provide clearer evidence for the intuition of our proposed DaM and HOG reconstruction mechanism, we calculate the distribution distances of the feature representations across different target domains. Following prior works [41, 1, 30], we compute the Jensen–Shannon ($JS$) divergence between two adjacent domains to represent the inter-domain divergence. If the inter-domain divergence is small, it indicates that the feature representation remains consistent and is less susceptible to cross-domain shifts [15]. For comparative experiments, we compute the inter-domain $JS$ divergence of the source model, using the DaM mechanism and our proposed method on the CIFAR10-to-CIFAR10C scenario. As illustrated in Fig. 4, when using the DaM mechanism, the inter-domain divergence is smaller than the source model on the vast majority of adjacent domains. Our method achieves the minimum inter-domain divergence on all fourteen adjacent domains. The observed pattern in inter-domain divergence suggests that DaM excels in extracting target domain knowledge but may exhibit limitations in robustness during continual adaptation. Meanwhile, after the HOG reconstruction of masked tokens, the model attains increased stability in cross-domain learning, mitigating the impact of domain shift erosion.

### v2

# Continual-MAE: Adaptive Distribution Masked Autoencoders for Continual Test-Time Adaptation

@@ Abstract
Continual Test-Time Adaptation (CTTA) is proposed to migrate a source pre-trained model to continually changing target distributions, addressing real-world dynamism. Existing CTTA methods mainly rely on entropy minimization or teacher-student pseudo-labeling schemes for knowledge extraction in unlabeled target domains. However, dynamic data distributions cause miscalibrated predictions and noisy pseudo-labels in existing self-supervised learning methods, hindering the effective mitigation of error accumulation and catastrophic forgetting problems during the continual adaptation process. To tackle these issues, we propose a continual self-supervised method, Adaptive Distribution Masked Autoencoders (ADMA), which enhances the extraction of target domain knowledge while mitigating the accumulation of distribution shifts. Specifically, we propose a Distribution-aware Masking (DaM) mechanism to adaptively sample masked positions, followed by establishing consistency constraints between the masked target samples and the original target samples. Additionally, for masked tokens, we utilize an efficient decoder to reconstruct a hand-crafted feature descriptor (e.g., Histograms of Oriented Gradients), leveraging its invariant properties to boost task-relevant representations. Through conducting extensive experiments on four widely recognized benchmarks, our proposed method attains state-of-the-art performance in both classification and segmentation CTTA tasks. Our project page: https://sites.google.com/view/continual-mae/home
@@ 1 Introduction
Deep Neural Networks (DNNs) have demonstrated impressive performance across various computer vision tasks, including image-level classification [20, 10, 34], dense prediction [44, 66, 55], and multi-model task [59, 58, 31], when the test data distribution closely aligns with the training data. However, this assumption is frequently challenged in real-world scenarios due to dynamic environments, with deployed models exhibiting insufficient generalization capabilities and performance degradation [47, 23]. Therefore, the problem of continual test-time adaptation (CTTA) has been introduced [52], aiming to adapt a source pre-trained model to continually changing target distributions. Due to privacy and practical considerations, during the continual adaptation process, access to source domain data is not permitted, and each target data can only be accessed once. While the CTTA showcases promising potential applications, it also increases the difficulty of transfer learning, which introduces catastrophic forgetting and error accumulation problems.
@@ 2.1 Continual Test-Time Adaptation
Test-time adaptation (TTA), also known as source-free domain adaptation [5, 29, 57, 62], is the process of adapting a source model to a target domain distribution without relying on any source domain data. Recent works have delved into techniques such as self-training and entropy regularization to fine-tune the source model [32, 6]. Tent [51] achieves this by updating the training parameters in batch normalization layers through entropy minimization. This approach has served as inspiration for subsequent research efforts in recent works [38, 61], which continue to investigate the robustness of normalization layers.And  [14] first attempts to reconstruct RGB images in the TTA task. Continual Test-Time Adaptation (CTTA) denotes a scenario where the target domain is dynamically changing, introducing additional challenges for conventional TTA methods. The initial approach is presented in [52], which employs a teacher-student framework to integrate bi-average pseudo labels and stochastic weight reset. Drawing from the insight that mean teacher predictions are often more robust than standard models [49], a series of mainstream methods [13, 33, 9] continue this approach to self-supervised learning in CTTA. Concurrently, existing methods also focus on applying entropy minimization to update normalization layers [50, 18] or a fraction of model parameters [48, 36]. However, due to continually changing environments, these self-training approaches are susceptible to miscalibrated predictions and noisy pseudo-labels, resulting in uncontrollable error accumulation.
@@ 2.2 Masked Image Modeling
Mask-reconstruction-based self-supervised learning has been successful in reducing the reliance on extensive labeled datasets in both Natural Language Processing (NLP) and Computer Vision (CV). The concept was first introduced by BERT [8] in NLP, where a portion of the input word tokens is randomly masked, and the model learns to reconstruct the vocabularies of these masked tokens. In the field of computer vision (CV), similar techniques have been applied in various works [3, 22, 56]. These methods involve randomly masking a significant percentage of input image patches. Specifically, BEiT [3] was the first to explore Masked Image Modeling (MIM) in vision transformers by reconstructing the vision dictionary derived from DALL-E [42, 43]. MAE [22] scaled up MIM to larger models and demonstrated that a simple pixel reconstruction loss can enhance the visual representations of pre-trained models. However, relying solely on low-level RGB signals in MAE is considered rudimentary and limited in unlocking the full potential of MAE in downstream vision tasks. Subsequently, approaches like MaskFeat [53], data2vec [2], MVP [54], and MILAN [24] have uncovered various high-level signals [65, 19], including pre-trained DINO features [39], hand-crafted features [7], momentum features [21], and multi-modality features [41]. Utilizing these high-level signals has been proven to be more effective in extracting contextual information.
@@ 3.1 Overview
Adaptive Distribution Masked Autoencoders Prevalent CTTA approaches primarily focus on applying entropy minimization to update the batch normalization layer [50, 18, 37] or a fraction of model parameters [48]. However, due to the dynamic environments, this self-training approach is susceptible to miscalibrated predictions, resulting in uncontrollable error accumulation. On the other hand, alternative mainstream approaches utilize the teacher-student pseudo-labeling in continual target domains. Nevertheless, the conventional mean teacher method [49] increases computational cost and produces noisy pseudo labels in dynamic environments, leading to the accumulation of distribution shifts. While methods such as [52, 33, 9] utilize Test-time augmentation methods to enhance the accuracy of pseudo labels, they may limit efficiency during the CTTA process. Different from prior continual self-supervised approaches, we lead the way in incorporating a masked autoencoder (MAE) to tackle the CTTA problem, abandoning the impact of miscalibrated predictions and the cumbersome teacher-student model. Our key insight lies in adopting the reconstruction scheme to effectively extract target domain knowledge while mitigating the accumulation of domain shift. The overall framework is illustrated in Figure 2. Specifically, We propose a Distribution-aware Masking (DaM) mechanism, detailed in Sec. 3.2, which adaptively places learnable masks on token embeddings with substantial distribution shifts. By establishing consistency constraints between the masked input and the original input, DaM significantly enhances the understanding of target domain knowledge and mitigates the challenge of error accumulation. For masked tokens, we adopt a linear decoder to reconstruct Histograms of Oriented Gradients, leveraging its invariant properties to acquire task-relevant representations while averting the introduction of target domain shifts, as elaborated in Sec. 3.3. This reconstruction method serves as a preventive measure, avoiding catastrophic forgetting in continual adaptation. We provide the intuitive explanation and justification in Sec. 5.
@@ 3.2 Distribution-aware Masking
To this end, as shown in Figure 2 (a), we introduce a Distribution-aware Masking (DaM) strategy, enabling a thoughtful selection of tokens to mask in dynamic environments. The key concept involves choosing tokens with substantial domain shifts for masking, ensuring that the preserved visible tokens exhibit relatively fewer domain shifts while providing reliable semantic knowledge through the model encoder. To quantify token-wise distribution shifts, we draw inspiration from [64, 63, 40, 45] and introduce a method for token-wise uncertainty estimation. Specifically, we employ the MC Dropout [11], enabling multiple forward propagations to obtain $m$ (e.g., $m=10$) sets of features for each token. Subsequently, we calculate the uncertainty value $\mathcal{U}(x)$ for a given token $x_{j}$, as formulated below:
@@ 4 Experiments
4.1 Experiments Details
Implementation Details. In our CTTA experiments, to ensure consistency and fair comparisons, we follow the implementation details as proposed in the prior CTTA works [52, 33]. For the classification CTTA tasks, we employ the ViT-base [10] as our backbone model. We resize the input images to 384$\times$384 for CIFAR10C, CIFAR100C benchmark, and 224$\times$224 resolution for ImageNet-C benchmark. In the case of segmentation CTTA task, we employ the Segformer-B5 [55] pre-trained on the Cityscapes dataset as our source model. We down-sample the original input images resolution from 1920$\times$1080 to 960$\times$540. We use Adam [26] with $(\beta_{1},\beta_{2})=(0.9,0.999)$ as the optimizer. Different learning rates are assigned for different CTTA tasks and backbone models, such as we use 1e-5 for ViT on CIFAR10C and CIFAR100C, 1e-3 for ViT on ImageNetC, and 3e-4 for Segformer on ACDC. In the masking process, following the approach in MIC [25], we use [MASK] tokens initialized with all zeros to replace 50% of tokens that are embedded as patches. These [MASK] tokens are shared learnable embeddings that indicate masked patches. The reconstruction decoder is a randomly initialized linear layer used to project the output tokens associated with masked patches to HOG features. The parameters of mask tokens and projection layer are optimized in parallel with other parameters. To achieve better results, we inject ViDA, as proposed by previous SOTA work [33], into the model and utilize our proposed self-supervised method to update.
4.2 Classification CTTA Tasks
CIFAR10-to-CIFAR10C & CIFAR100-to-CIFAR100C. The source model is trained through supervised learning on the CIFAR10 or CIFAR100 dataset. During testing, we apply CTTA to the CIFAR10C or CIFAR100C dataset, which contains fifteen corruption types continuously fed into the model in a specific order. In the CIFAR10-to-CIFAR10C scenario, as shown in Table 1, the average classification error of the source model reaches 28.2% when directly testing on CIFAR10C. However, our method significantly reduces the error to 12.6%. Compared to other continual self-supervised methods, our approach outperforms them by 10.9% and 12.0% compared to the previous entropy minimization method (TENT [51]) and the teacher-student method (CoTTA [52]), demonstrating significant potential in addressing the CTTA problem. To be mentioned, our Adaptive Distribution Masked Autoencoders (ADMA) demonstrate outstanding performance across 12 out of the 15 corruption types, validating the robustness of our method in the continual adaptation process. Note that, for TENT, we implement entropy minimization to update the Layer Normalization layers in the transformer instead of BN.
We expand our evaluation to the CIFAR100-to-CIFAR100C scenario, as depicted in Table 2, encompassing a more extensive range of categories within each domain. Our approach outperforms the previous entropy minimization and teacher-student pseudo-labeling methods by 5.7% and 8.4%, respectively, exhibiting a consistent trend with the aforementioned CTTA experiments. Therefore, the results validate the universality of our method, unaffected by the number of categories, and it can efficiently mitigate error accumulation and catastrophic forgetting problems.
@@ 4.4 Ablation Study
Error rate 49.2 48.5 48.1 43.6
Target SIFT Sobel Laplacian ORB
Error rate 50.4 48.2 48.7 50.0
Table 6: Average error rate(%) for ImageNet-to-ImageNet-C. RGB, CLIP, Mean-Teacher, SIFT [35], ORB, edge detectors(Sobel and Laplacian), and HOG are different reconstruction target.
@@
Reconstruction target selection. Another set of ablation experiments aims to assess the impact of the reconstruction target. This includes the original RGB pixel, high-level CLIP features, Mean-Teacher features, as well as SIFT, ORB, Sobel, Laplacian, and HOG features. Since the CIFAR datasets have an input size of $32\times 32$ with limited RGB pixel information, these ablation experiments are conducted in the context of the ImageNet-to-ImageNet-C. As shown in Table 6, our introduced HOG reconstruction outperforms other reconstruction targets by a significant margin. Specifically, compared with classical image pixel reconstruction, our method achieves a 5.6% improvement, demonstrating that reconstructing the RGB signal of the target domain introduces inherent domain shifts. In contrast to reconstructing the CLIP feature, our method outperforms it by 4.9%, while also significantly reducing computational costs as it does not require an additional model. The results validate that features directly extracted by the large-scale model still introduce domain shifts. Lastly, while reconstructing features of the Mean-Teacher model is slightly better (by 0.4%) than reconstructing CLIP features, it is still 4.5% lower than our method. In conclusion, reconstructing HOG and leveraging its invariant properties can boost task-relevant representations and avoid domain shift accumulation in CTTA.
@@ 5 Discussion and Justification
Inter-Domain Divergence. To provide clearer evidence for the intuition of our proposed DaM and HOG reconstruction mechanism, we calculate the distribution distances of the feature representations across different target domains. Following prior works [46, 1, 33], we compute the Jensen–Shannon ($JS$) divergence between two adjacent domains to represent the inter-domain divergence. If the inter-domain divergence is small, it indicates that the feature representation remains consistent and is less susceptible to cross-domain shifts [16]. For comparative experiments, we compute the inter-domain $JS$ divergence of the source model, using the DaM mechanism and our proposed method on the CIFAR10-to-CIFAR10C. As illustrated in Figure 4, when using the DaM mechanism, the inter-domain divergence is smaller than the source model on the vast majority of adjacent domains. Our method achieves the minimum inter-domain divergence on all fourteen adjacent domains. The observed pattern in inter-domain divergence suggests that DaM excels in extracting target domain knowledge but may exhibit limitations in robustness during continual adaptation. Meanwhile, after the HOG reconstruction of masked tokens, the model attains increased stability in cross-domain learning, mitigating the impact of domain shift erosion.